OpenStudy (anonymous):

Use the geometric mean to find the 7th term in a geometric sequence if the 6th term is 75 and the 8th term is 48.

OpenStudy (anonymous):

let 7 th term be 'x'.Then 75,x,48 are in GP. So,in GP x^2 =(75)(48) x^2 = (3600) x = square root of (3600) x = 60

OpenStudy (jskhupmang):

@PrincessKhiayla weren't you talking about squence and series?

OpenStudy (anonymous):

No, the answer was 60 @jskhupmang

OpenStudy (jskhupmang):

how did you know the answer?

OpenStudy (jskhupmang):

opps nvm

OpenStudy (jskhupmang):

so he was correct?